OpenStudy (anonymous):

What is the integral of (x^2+1/ x) dx from 1 to e?

OpenStudy (anonymous):

\[\int\limits_{1}^{e} (x^2+1) / x\]

OpenStudy (dumbcow):

split it into 2 fractions x^2/x = x

OpenStudy (dumbcow):

integral of 1/x is ln(x)

OpenStudy (anonymous):

so it becomes x^2/2 + ln(x)

OpenStudy (dumbcow):

correct, then evaluate F(e) -F(1) where F = x^2/2 +ln(x)

OpenStudy (anonymous):

i get (e^2+1/2) - (1/2 - 0) . Did i evaluate correctly?

OpenStudy (dumbcow):

the first part should be e^2/2 + 1

OpenStudy (anonymous):

Oh yeah thats corect. So when combined is (e^2 - 1 / 2)

OpenStudy (anonymous):

correction that subtraction should be addition

OpenStudy (anonymous):

so the final answer is (e^2+1/2)

OpenStudy (dumbcow):

yes ...if that is (e^2 +1)/2

OpenStudy (anonymous):

right thats what i have. Thank you so much for explaining. =) Im so bad with integrals.

OpenStudy (dumbcow):

no problem

OpenStudy (anonymous):

Do u mind helping me with another problem. im posting it now.